# Cartan–Thullen theorem and Levi problem in context of generalised convexity

### v2

@@ 3. Gelfand transform
    $\Phi$ is a homeomorphism of $\Omega$ and $\Phi(\Omega)$,
@@
Proof.
We shall first show that $\Phi$ is homeomorphism onto $\Phi(\Omega)$. As $\mathcal{A}$ separates points of $\Omega$ and consists of continuous functions, $\Phi$ is injective and continuous. Therefore, for any compact $K\subset\Omega$, the restriction of $\Phi$ to $K$ is a homeomorphism onto the corresponding image. As $\Omega$ is exhaustible by compactae, cf. Remark 2.3, there exist compact sets $(K_{i})_{i=1}^{\infty}$ such that
@@
This proves (iii). Point (v) is proven similarly. ∎
@@
If $\Omega$ is exhaustible by compactae and separable, $\mathcal{A}$ is closed in $\tau_{\mathcal{K}}$ and separates points of $\Omega$, then $\Omega$ is metrisable. By the Urysohn theorem it suffices to show that $\Omega$ has a countable basis of neighbourhoods and that it is completely regular.
As $\Omega$ is a countable union of open sets which are metrisable, thus it admits itself a countable basis of neighbourhoods. Lemma 3.1 shows that the topology on $\Omega$ is the weak topology induced by $\mathcal{A}$. Thus, it is completely regular.
4. $\mathcal{F}$-convex sets
@@
Proof.
Any element of $\mathcal{F}$ is a supremum of elements of $\mathcal{A}$, by Lemma 2.8. The claim follows. ∎
@@
An application of Lemma 4.2 completes the proof. ∎
@@
We shall say that $S\subset\Omega$ is $\mathcal{F}$-convex whenever
$S=\mathrm{Conv}_{\mathcal{F}}S.$
Here $G$ is the complete lattice cone of convex, lower semi-continuous functions on $\mathcal{A}^{*}$ generated by evaluations on elements of $\mathcal{A}$.
@@
Note that above $\mathrm{Conv}_{G}\Phi(S)$ denotes the usual convex hull of a set $\Phi(S)\subset\mathcal{A}^{*}$. Moreover, if $S\subset\Omega$ is closed, then it is $\mathcal{F}$-convex if and only if
$S=\mathrm{clConv}_{\mathcal{F}}S.$
This is to say, Definition 4.4 and Definition 4.1 are consistent.
@@
Proof.
The claim follows by Defnition 4.4, Lemma 3.1, (iv) and an analogous claim for convex, lower semi-continuous functions on $\mathcal{A}^{*}$. ∎
@@
That is $\omega$ belongs to $\mathrm{clConv}_{\mathcal{F}}K$. It is trivial that (iii) implies (i). ∎
@@
Let us note that the above does not imply that for any $K\subset\Omega$ and any $C\geq 1$ the set $\mathrm{clConv}_{\mathcal{F}}K$ is equal to
@@
consists of the elements of the affine hull of a compact set $K$. Since $K$ had non-empty interior, its affine hull contains $\Omega$. Thus, it is not equal to $K$.
@@
This shows that $\omega\in\mathrm{clConv}_{\mathcal{F}}K$ and concludes the proof. ∎
@@
Let $\mathcal{F}$ be a convex cone of of continuous functions on $\Omega$ that contains constants. Then for any compact set $K\subset\Omega$,
$\{\omega\in\Omega\mid\lvert f(\omega)\rvert\leq\sup\lvert f\rvert(K)\text{ forall }f\in\mathcal{F}\}=\mathrm{clConv}_{\mathcal{F}}K.$
Indeed, all functions in $\mathcal{F}$ are continuous, thus if $f\in\mathcal{F}$, then $f+\inf f(K)\in\mathcal{F}$. If $\omega$ belongs to $\mathrm{clConv}_{\mathcal{F}}K$, then for all $f\in\mathcal{F}$,
$f(\omega)+\inf f(K)\leq\sup f(K)+\inf f(K),$
so that $\omega$ belongs also to the set on the right-hand side of the above equality. The converse inclusion is trivial.
5. Proofs
5.1. Proof of Theorem 1.11
@@ 5. Proofs
Let $\Omega$ be a topological space. Let $\mathcal{A}$ be a linear space of continuous functions on $\Omega$ that contains constants and let $\mathcal{F}$ be the complete lattice cone generated by $\mathcal{A}$. Suppose that $\Omega$ is exhaustible by compactae and that for any compact set $K\subset\Omega$ and $C\geq 1$, the set
@@
for some symmetric family $(a_{\alpha})_{\alpha\in A}$ of functions in $\mathcal{A}$. Moreover for any compact set $K$, $p$ is a maximum of a finite number of functions of the family.
Proof.
We shall first choose a sequence of compactae $(K_{i})_{i=1}^{\infty}$, such that for each $i=1,2\dotsc,$
@@
Thus, the sequence of the sets is constructed.
We shall $\mathcal{F}$ denote the complete lattice cone generated by $\mathcal{A}$.
@@
$p_{i}=\max\Big{\{}\frac{\lvert a_{l}\rvert}{2^{i}\lVert a_{l}\rVert_{\mathcal{C}(K_{i})}}\mid l=1,\dotsc,k_{i}\Big{\}}\vee 0.$
@@
This completes the proof. ∎
Proof of Theorem 1.11.
By Lemma 5.1 and Remark 2.3 it follows that (i) implies (ii). Clearly (ii) implies (iii) which in trun implies (iv).
We shall $\mathcal{F}$ denote the complete lattice cone generated by $\mathcal{A}$.
@@
Then $K\subset p^{-1}([0,t])$ implies that for $\alpha\in A$, $\sup\lvert a_{\alpha}\rvert(K)\leq t$. Let $\omega\in S$. Then we see
@@
Let us now show that (ii) implies (iv).
Recall that $p$ on any compact is a maximum of a finite family of functions in $\mathcal{A}$. Let
@@
This is to say, (iv) is satisfied. If $\mathcal{A}$ consists of real-parts of a complex algebra, Proposition 4.9 shows that (i) is equivalent to compactness of all sets $\mathrm{clConv}_{\mathcal{A}}K$, for compact $K\subset\Omega$. Since for any such $K$ there is some $i=1,2,\dotsc$ such that $K\subset P_{i}$, and these $P_{i}$ are $\mathcal{A}$-convex, we see that (i) follows.
We shall now show that (i) is equivalent to $\Omega$ being complete with respect to $\mathrm{cl}\mathcal{A}$, where the closure is taken with respect to compact-open topology $\tau_{\mathcal{K}}$.
Since $\Omega$ is separable, it is $\Omega$ is metrisable, see Remark 3.2. Let us pick a Cauchy sequence $(\omega_{n})_{n=1}^{\infty}$ of elements of $\Omega$. The sequence induces a sequence $(\Phi(\omega_{n}))_{n=1}^{\infty}$ of linear, continuous functionals on $\mathcal{A}$, which is pointwise bounded. Therefore, by the Banach–Steinhaus uniform boundedness principle we can find a compact set $K\subset\Omega$ and a constant $C>0$ such that for all $a\in\mathcal{A}$ and all $n=1,2,\dotsc$
@@
which is compact by (i). It follows that $(\omega_{n})_{n=1}^{\infty}$ converges. We have shown that $\Omega$ is complete with respect to $\mathcal{A}$.
Suppose now conversely that $\Omega$ is complete. Recall that by Lemma 3.1, (i), the map $\Phi\colon\Omega\to\mathcal{A}^{*}$ is a homeomorphism onto its image, when $\mathcal{A}^{*}$ is equipped with the weak* topology. Let $K\subset\Omega$ and let $C\in\mathbb{R}$. Then the set
@@
Let us assume that (i) is satisfied. Suppose that $\Omega$ is locally compact, separable, that $\mathcal{A}$ consists of real-parts of a complex algebra. Remark 3.2 shows that $\Omega$ is metrisable. We shall show that $\Omega$ is an $\mathrm{cl}\mathcal{A}$-space.
@@
Let $\bar{\Omega}$ denote the completion of $\Omega$. Then $\bar{\Omega}$ is separable as a metrisable space with a separable dense subset $\Omega$. It admits a countable basis $(U_{i})_{i=1}^{\infty}$ of neighbourhoods of points in $\bar{\Omega}\setminus\Omega$. Let us take a function $k\colon\mathbb{N}\to\mathbb{N}$, with the property that each element of $\mathbb{N}$ appears infinitely often in its image. As the sets $(K_{j})_{j=1}^{\infty}$ are compact, for any $j=1,2,\dotsc$ there exists an element $\omega_{j}\in U_{k(j)}\cap\Omega\setminus K_{j}$. Relabelling the sets $(K_{j})_{j=1}^{\infty}$ if necessary, we may assume that $\omega_{j}\in K_{j+1}$. By Proposition 4.9 there exists $a_{j}\in\mathcal{A}$ such that
@@
(5.1) $a(\omega_{j})>a_{j}(\omega_{j})-\sum_{i=1}^{j-1}\lvert a_{i}\rvert(\omega_{j})-\sum_{i=j+1}^{\infty}\lvert a_{i}\rvert(\omega_{j})\geq j-2^{-j}.$
We shall now show that $a$ verifies the fact that $\Omega$ is an $\mathrm{cl}\mathcal{A}$-space. We shall show that it does not extend to a continuous function on $\Omega\cup\{\omega\}$ for any $\omega\in\bar{\Omega}\setminus\Omega$. Suppose on the contrary, that it does extend to $\omega\in\bar{\Omega}\setminus\Omega$. By (5.1), $a$ is takes arbitrary large values on any neighbourhood of $\omega$. If it extended continuously to $\omega$, it would have to be infinite on $\omega$.
As already observed, the fact that $\Omega$ is an $\mathrm{cl}\mathcal{A}$-space immediately implies that $\Omega$ is complete with respect to $\mathrm{cl}\mathcal{A}$. ∎
@@
The Mazur theorem and Lemma 3.1 imply that $\mathrm{clConv}_{G}\Phi(K)$ is compact. The proof of Theorem 1.11 shows that $\Phi(\Omega)$ is closed in the weak* topology on $\mathcal{A}^{*}$, so that
@@
is a compact subset of $\Phi(\Omega)$. As $\Phi$ is a homeomorphism onto its image, so that $\mathrm{clConv}_{\mathcal{F}}K$ is compact subset of $\Omega$.
@@
Let us recall the following theorem of Krein and Šmulian, [23]. The original version was formulated for normed spaces. We shall be interested in a version for Fréchet spaces, see [37, Theorem 3., p. 190, Definition 1., p. 184]. We cite it below.
@@
If $\Phi(\Omega)\subset\mathcal{A}^{*}$ was convex, then it would follow from the Krein–Šmulian theorem. Indeed, in that case $\Phi(\Omega)$ is weakly* closed if and only if for any $C\geq 1$ and compact set $K\subset\Omega$ the set
@@
The crucial observation in the proof of Theorem 1.11 is that $\Omega$ is metrisable, so that to prove its completeness it suffices to investigate behaviour of Cauchy sequences. Now, if $(\omega_{n})_{n=1}^{\infty}$ is a Cauchy sequence, then $(\Phi(\omega_{n}))_{n=1}^{\infty}$ is a pointwise bounded sequence of linear functionals on $\mathcal{A}$. This would not be necessarily true, if we were to consider general nets of functionals on $\mathcal{A}$. Now, as $(\Phi(\omega_{n}))_{n=1}^{\infty}$ is pointwise bounded, we are free to employ the Banach–Steinhaus uniform boundedness principle and in conclusion infer completeness.
If $\mathcal{A}$ is consists of real-parts of a closed complex algebra, then there is an alternative proof of the fact that (i) of Theorem 1.11 implies completeness of $\Omega$ with respect to $\mathrm{cl}\mathcal{A}$, that is analogous to the proof [17, Theorem 7, p. 65]. It relies on the following lemma.
@@
We shall proceed inductively. Since $\Omega$ is exhaustible by compactae, there is a sequence of compact sets $(L_{i})_{i=1}^{\infty}$ for which
(5.2) $\bigcup_{i=1}^{\infty}L_{i}=\Omega,L_{i}\subset\mathrm{int}L_{i+1}\text{ for }i=1,2,\dotsc.$
Suppose the functions $a_{1},\dotsc,a_{k}\in\mathcal{A}$, elements $n_{1},\dotsc,n_{k}$ and sets $K_{1},\dotsc,K_{k}$ are chosen. There is an index $j_{k}\geq k$ such that $K_{k}\cup\{\omega_{n_{1}},\dotsc,\omega_{n_{k}}\}\subset\mathrm{int}L_{j_{k}}$. By the assumption, the set
@@
Thus, there also exists $b_{k+1}\in\mathcal{A}$ for which
$\lvert b_{k+1}\rvert(\omega_{n_{k+1}})>\sup\lvert b_{k+1}\rvert(K_{k+1}).$
Normalising, taking power sufficiently high power and then real-part of a multiple of $b_{k+1}$ yields a function $a_{k+1}$ that we were after. Since $j_{k}\geq k$, we see that
@@
We define now
$a=\sum_{i=1}^{\infty}a_{i}.$
Since
$\sup\lvert a_{i}\rvert(K_{j})\leq 2^{-i}\text{ for }i\geq j,$
we see that the series converges in $\tau_{\mathcal{K}}$, and therefore $a\in\mathrm{cl}\mathcal{A}$. However, for $j=1,2,\dotsc$ we have
$a(\omega_{n_{j}})\geq a_{j}(\omega_{n_{j}})-\sum_{i=1}^{j-1}\lvert a_{i}\rvert(\omega_{n_{j}})-\sum_{i=j+1}^{\infty}\lvert a_{i}\rvert(\omega_{n_{j}})\geq j-2^{-j},$
as $\omega_{n_{j}}\in K_{i}$ for $i\geq j+1$, so that
$\lvert a_{i}\rvert(\omega_{n_{j}})\leq 2^{-i}.$
This shows that
$\lim_{j\to\infty}a(\omega_{j})=\infty.$
Remark 2.6 and Remark 3.2 show that $\Omega$ is metrisable. To prove that $\Omega$ is complete with respect to $\mathcal{A}$, it suffices to show that any Cauchy sequence $(\omega_{n})_{n=1}^{\infty}$ is convergent. Lemma 5.6 shows that any Cauchy sequence has an accumulation point, and therefore it is convergent in $\Omega$. Thus Lemma 5.6 proves that (i) of Theorem 1.11 implies completeness of $\Omega$.
5.2. Proof of Theorem 1.18
@@
Proof.
As $K_{1},K_{2}$ are compact and $\mathcal{F}$-convex we see that for $i=1,2$
@@
(5.3) $\Big{\{}\omega\in\Omega\mid f(\omega)>\sup f(K_{2})\Big{\}},\text{ for all }f\in\mathcal{F},$
@@
(5.4) $\Big{\{}\omega\in K_{4}\mid\frac{f_{i}(\omega)-\sup f_{i}(K_{1})}{\sup f_{i}(K_{2})-\sup f_{i}(K_{1})}\leq 1\text{ for }i=1,\dotsc,k\Big{\}}\subset\mathrm{int}K_{3}.$
@@
(5.5) $p=\max\Big{\{}\frac{f_{i}-\sup f_{i}(K_{1})}{\sup f_{i}(K_{2})-\sup f_{i}(K_{1})}\mid i=1,\dotsc,k\Big{\}}\vee 0$
satisfies our requirements. ∎
@@
(5.7) $p\geq i\text{ on }\Omega\setminus\mathrm{int}K_{i+2}.$
@@
By continuity of $p$ we see that preimages of compactae are compact, i.e., $p$ is proper.
This completes the proof. ∎
Proof of Theorem 1.18.

### v3

@@ 3 Gelfand transform
    $\Phi$ is a homeomorphism between $\Omega$ and $\Phi(\Omega)$,
@@
Proof 3.2.
We shall first show that $\Phi$ is homeomorphism onto $\Phi(\Omega)$. As $\mathcal{A}$ separates points of $\Omega$ and consists of continuous functions, $\Phi$ is injective and continuous. Therefore, for any compact $K\subset\Omega$, the restriction of $\Phi$ to $K$ is a homeomorphism onto the corresponding image. As $\Omega$ is exhaustible by compacta, cf. Remark 2.3, there exist compact sets $(K_{i})_{i=1}^{\infty}$ such that
@@
This proves (iii). Point (v) is proven similarly.
@@
If $\Omega$ is exhaustible by compacta and separable, $\mathcal{A}$ is closed in $\tau_{\mathcal{K}}$ and separates points of $\Omega$, then $\Omega$ is metrisable. By the Urysohn theorem it suffices to show that $\Omega$ has a countable basis of neighbourhoods and that it is completely regular.
As $\Omega$ is a countable union of open sets which are metrisable, hence it admits a countable basis of neighbourhoods. Lemma 3.1 shows that the topology on $\Omega$ is the weak topology induced by $\mathcal{A}$. Thus, it is completely regular.
4 $\mathcal{F}$-convex sets
@@
Proof 4.3.
Any element of $\mathcal{F}$ is a supremum of elements of $\mathcal{A}$, by Lemma 2.8. The claim follows.
@@
An application of Lemma 4.2 completes the proof.
@@
We shall say that $S\subset\Omega$ is $\mathcal{F}$-convex whenever $S=\mathrm{Conv}_{\mathcal{F}}S$. Here $G$ is the complete lattice cone of convex, lower semi-continuous functions on $\mathcal{A}^{*}$ generated by evaluations on elements of $\mathcal{A}$.
@@
Note that above $\mathrm{Conv}_{G}\Phi(S)$ denotes the usual convex hull of a set $\Phi(S)\subset\mathcal{A}^{*}$. Moreover, if $S\subset\Omega$ is closed, then it is $\mathcal{F}$-convex if and only if $S=\mathrm{clConv}_{\mathcal{F}}S$. This is to say, Definition 4.6 and Definition 4.1 are consistent.
@@
Proof 4.9.
The claim follows by Defnition 4.6, Lemma 3.1, (iv) and an analogous claim for convex, lower semi-continuous functions on $\mathcal{A}^{*}$.
@@
That is, $\omega$ belongs to $\mathrm{clConv}_{\mathcal{F}}K$. It is trivial that (iii) implies (i).
@@
Note that the above does not imply that for any $K\subset\Omega$ and any $C\geq 1$ the set $\mathrm{clConv}_{\mathcal{F}}K$ is equal to
@@
consists of the elements of the affine hull of the compact set $K$. Since $K$ had non-empty interior, its affine hull contains $\Omega$. Thus, it is not equal to $K$.
@@
This shows that $\omega\in\mathrm{clConv}_{\mathcal{F}}K$ and concludes the proof.
@@
Let $\mathcal{A}$ be a linear space of continuous functions on $\Omega$ that contains constants. Then for any compact set $K\subset\Omega$,
$\{\omega\in\Omega\mid\lvert a(\omega)\rvert\leq\sup\lvert a\rvert(K)\text{ forall }a\in\mathcal{A}\}=\mathrm{clConv}_{\mathcal{A}}K.$ (4.1)
Indeed, all functions in $\mathcal{A}$ are continuous, thus if $a\in\mathcal{A}$, then $a+\inf a(K)\in\mathcal{A}$. If $\omega$ belongs to the set on the left-hand side of (4.1), then for all $a\in\mathcal{A}$,
$a(\omega)+\inf a(K)\leq\sup a(K)+\inf a(K),$
so that $\omega$ belongs also to $\mathrm{clConv}_{\mathcal{A}}(K)$. The converse inclusion follows by Lemma 4.2.
5 Proofs
5.1 Proof of Theorem 1.11
@@ 5 Proofs
Let $\Omega$ be a topological space. Let $\mathcal{A}$ be a linear space of continuous functions on $\Omega$ that contains constants and let $\mathcal{F}$ be the complete lattice cone generated by $\mathcal{A}$. Suppose that $\Omega$ is exhaustible by compacta and that for any compact set $K\subset\Omega$ and $C\geq 1$, the set
@@
for some symmetric family $(a_{\alpha})_{\alpha\in A}$ of functions in $\mathcal{A}$. Moreover for any compact set $K$, $p$ on $K$ is a maximum of a finite number of the functions in the family.
The lemma above is new, up to the author’s knowledge, and the main new ingredient is the exploitation of the compactness of the sets
for arbitrary $C\geq 1$, which in the setting of complex analysis is not widely studied.
Proof 5.2.
We shall first choose a sequence of compacta $(K_{i})_{i=1}^{\infty}$, such that for each $i=1,2\dotsc,$
@@
Thus, the required sequence of the sets is constructed.
@@
$p_{i}=\max\Bigg{\{}\frac{\lvert a_{l}\rvert}{2^{i}\lVert a_{l}\rVert_{\mathcal{C}(K_{i})}}\mid l=1,\dotsc,k_{i}\Bigg{\}}\vee 0.$
@@
This completes the proof.
Proof 5.3 (Proof of Theorem 1.11).
By Lemma 5.1 and Remark 2.3 it follows that (i) implies (ii). Clearly, (ii) implies (iii) which in turn implies (iv).
Let $\mathcal{F}$ denote the complete lattice cone generated by $\mathcal{A}$.
@@
Then $K\subset p^{-1}([0,t])$ implies that for $\alpha\in A$, $\sup\lvert a_{\alpha}\rvert(K)\leq t$. Let $\omega\in S$. Then
@@
Let us now show that (ii) implies existence of a family $(P_{i})_{i=1}^{\infty}$ of compact, symmetric $\mathcal{A}$-polytopes that exhaust $\Omega$ and such that $P_{i}\subset\mathrm{int}P_{i+1}$ for $i=1,2,\dotsc$.
By the assumption, $p$ on any compact set is a maximum of a finite family of functions in $\mathcal{A}$. Let
@@
This shows the existence of the requested $\mathcal{A}$-polytopes.
If $\mathcal{A}$ consists of real-parts of a complex algebra, Proposition 4.13 shows that (i) is equivalent to compactness of all sets $\mathrm{clConv}_{\mathcal{A}}K$, for compact $K\subset\Omega$. Since for any such $K$ there is some $i=1,2,\dotsc$ such that $K\subset P_{i}$, and these $P_{i}$ are $\mathcal{A}$-convex, we see that (i) follows.
Under the assumption that $\Omega$ is separable, we shall show that (i) is equivalent to $\Omega$ being complete with respect to $\mathrm{cl}\mathcal{A}$, where the closure is taken with respect to compact-open topology $\tau_{\mathcal{K}}$.
By Remark 3.3 $\Omega$ is metrisable. Therefore, to prove that $\Omega$ is complete with respect to $\mathrm{cl}\mathcal{A}$, it suffices to consider Cauchy sequences instead of general Cauchy nets. Let us pick a Cauchy sequence $(\omega_{n})_{n=1}^{\infty}$ of elements of $\Omega$. The sequence induces a sequence $(\Phi(\omega_{n}))_{n=1}^{\infty}$ of linear, continuous functionals on $\mathrm{cl}\mathcal{A}$, which is pointwise bounded. Therefore, by the Banach–Steinhaus uniform boundedness principle we can find a compact set $K\subset\Omega$ and a constant $C>0$ such that for all $a\in\mathcal{A}$ and all $n=1,2,\dotsc$
@@
which is compact by (i). It follows that $(\omega_{n})_{n=1}^{\infty}$ converges. We have shown that $\Omega$ is complete with respect to $\mathrm{cl}\mathcal{A}$.
Suppose now conversely that $\Omega$ is complete with respect to $\mathrm{cl}\mathcal{A}$. Recall that by Lemma 3.1, (i), the map $\Phi\colon\Omega\to\mathcal{A}^{*}$ is a homeomorphism onto its image, when $\mathcal{A}^{*}$ is equipped with the weak* topology. Let $K\subset\Omega$ and let $C\in\mathbb{R}$. Then the set
@@
Let us assume that (i) is satisfied. Suppose that $\Omega$ is separable and that $\mathcal{A}$ consists of real-parts of a complex algebra. Remark 3.3 shows that $\Omega$ is metrisable. Let $d$ be a metric on $\Omega$ that yields topology of $\Omega$. We shall show that $\Omega$ is an $(\mathrm{cl}\mathcal{A},d)$-space.
@@
Let $\bar{\Omega}$ denote the completion of $\Omega$ with respect to $d$. Then $\bar{\Omega}$ is separable as a metrisable space with a separable dense subset $\Omega$. It admits a countable basis $(U_{i})_{i=1}^{\infty}$ of neighbourhoods of points in $\bar{\Omega}\setminus\Omega$. Let us take a function $k\colon\mathbb{N}\to\mathbb{N}$, with the property that each element of $\mathbb{N}$ appears infinitely often in its image. As the sets $(K_{j})_{j=1}^{\infty}$ are compact, for any $j=1,2,\dotsc$ there exists an element $\omega_{j}\in U_{k(j)}\cap\Omega\setminus K_{j}$. Relabelling the sets $(K_{j})_{j=1}^{\infty}$ if necessary, we may assume that $\omega_{j}\in K_{j+1}$. By Proposition 4.13 there exists $a_{j}\in\mathcal{A}$ such that
@@
$a(\omega_{j})>a_{j}(\omega_{j})-\sum_{i=1}^{j-1}\lvert a_{i}\rvert(\omega_{j})-\sum_{i=j+1}^{\infty}\lvert a_{i}\rvert(\omega_{j})\geq j-2^{-j}.$ (5.1)
We shall now show that $a$ verifies the fact that $\Omega$ is an $(\mathrm{cl}\mathcal{A},d)$-space. Let $(\omega_{\alpha})_{\alpha\in A}$ be a Cauchy net with respect to $d$ that does not converge to an element in $\Omega$, so that its limit $\omega\in\bar{\Omega}\setminus\Omega$. By (5.1) we see that $\lim_{\alpha\in A}a(\omega_{\alpha})=\infty$, as $(a(\omega_{\alpha}))_{\alpha\in A}$ is unbounded on any neighbourhood of $\omega$.
Suppose now that $\Omega$ is an $(\mathrm{cl}\mathcal{A},d)$-space for any metric $d$ that yields the topology of $\Omega$. We shall show that $\Omega$ is complete with respect to $\mathrm{cl}\mathcal{A}$. Remark 2.6 shows that $\mathrm{cl}\mathcal{A}$ is separable in $\tau_{\mathcal{K}}$. Let $(a_{i})_{i=1}^{\infty}$ be a dense and countable subset of $\mathrm{cl}\mathcal{A}$. Define for $\omega_{1},\omega_{2}\in\Omega$
$d(\omega_{1},\omega_{2})=\sum_{i=1}^{\infty}\frac{1}{2^{i}}\frac{\lvert a_{i}(\omega_{1})-a_{i}(\omega_{2})\rvert}{\lvert a_{i}(\omega_{1})-a_{i}(\omega_{2})\rvert+1}.$ (5.2)
Then $d\colon\Omega\times\Omega\to\mathbb{R}$ is a metric on $\Omega$. Since $\Omega$ is exhaustible by compacta, we may use an argument similar to the one in Lemma 3.1 to show that above defined $d$ metrises the topology of $\Omega$.
Pick now any Cauchy net $(\omega_{\alpha})_{\alpha\in A}$ with respect to $\mathrm{cl}\mathcal{A}$. It is also a Cauchy net with respect to $d$, so that we may take its equivalence class $\omega\in\bar{\Omega}$. Suppose that $\omega$ is not equivalent to a point in $\Omega$. By the assumption, there exists $a\in\mathrm{cl}\mathcal{A}$ for which $\lim_{\alpha\in A}a(\omega_{\alpha})=\infty$. This is impossible, as the sequence is a Cauchy sequence with respect to $\mathrm{cl}\mathcal{A}$, so that $(a(\omega_{\alpha}))_{\alpha\in A}$ converges. Thus, $\Omega$ is complete with respect to $\mathrm{cl}\mathcal{A}$.
@@
The Mazur theorem and Lemma 3.1 imply that $\mathrm{clConv}_{G}\Phi(K)$ is compact. The proof of Theorem 1.11 shows that if $\Omega$ is complete, then $\Phi(\Omega)$ is closed in the weak* topology on $\mathcal{A}^{*}$. Therefore
@@
is a compact subset of $\Phi(\Omega)$. As $\Phi$ is a homeomorphism onto its image, we see that $\mathrm{clConv}_{\mathcal{F}}K$ is compact subset of $\Omega$.
@@
Let us recall the following theorem of Krein and Šmulian, [26]. The original version was formulated for normed spaces. We shall be interested in a version for Fréchet spaces, see [42, Theorem 3., p. 190, Definition 1., p. 184]. We cite it below.
@@
If $\Phi(\Omega)\subset\mathcal{A}^{*}$ was convex, then the equivalence would follow from the Krein–Šmulian theorem. Indeed, in that case $\Phi(\Omega)$ is weakly* closed if and only if for any $C\geq 1$ and compact set $K\subset\Omega$ the set
@@
The crucial observation in the proof of Theorem 1.11 is that $\Omega$ is metrisable, so that to prove its completeness it suffices to investigate behaviour of Cauchy sequences. Now, if $(\omega_{n})_{n=1}^{\infty}$ is a Cauchy sequence, then $(\Phi(\omega_{n}))_{n=1}^{\infty}$ is a pointwise bounded sequence of linear functionals on $\mathcal{A}$. This would not be necessarily true, if we were to consider general nets of functionals on $\mathcal{A}$. Now, as $(\Phi(\omega_{n}))_{n=1}^{\infty}$ is pointwise bounded, we are free to employ the Banach–Steinhaus uniform boundedness principle and infer completeness.
If $\mathcal{A}$ is consists of real-parts of a closed complex algebra, then there is an alternative proof of the fact that (i) of Theorem 1.11 implies completeness of $\Omega$ with respect to $\mathrm{cl}\mathcal{A}$, that is analogous to the proof [20, Theorem 7, p. 65]. It relies on the following lemma.
@@
We shall proceed inductively. Since $\Omega$ is exhaustible by compacta, there is a sequence of compact sets $(L_{i})_{i=1}^{\infty}$ for which
$\bigcup_{i=1}^{\infty}L_{i}=\Omega,L_{i}\subset\mathrm{int}L_{i+1}\text{ for }i=1,2,\dotsc.$ (5.3)
Suppose the functions $a_{1},\dotsc,a_{k}\in\mathcal{A}$, elements $n_{1},\dotsc,n_{k}$ and sets $K_{1},\dotsc,K_{k}$ are chosen. There is an index $j_{k}\geq k$ such that $K_{k}\cup\{\omega_{n_{1}},\dotsc,\omega_{n_{k}}\}\subset\mathrm{int}L_{j_{k}}$. By the assumption and Proposition 4.13, the set
@@
Thus, there also exists $b_{k+1}\in\mathcal{A}$ for which $\lvert b_{k+1}\rvert(\omega_{n_{k+1}})>\sup\lvert b_{k+1}\rvert(K_{k+1})$. Normalising, taking power sufficiently high power and then real-part of a multiple of $b_{k+1}$ yields a function $a_{k+1}$ that we were after. Since $j_{k}\geq k$, we see that
@@
We define now $a=\sum_{i=1}^{\infty}a_{i}$. Since $\sup\lvert a_{i}\rvert(K_{j})\leq 2^{-i}$ for $i\geq j$, we see that the series converges in $\tau_{\mathcal{K}}$, and therefore $a\in\mathrm{cl}\mathcal{A}$. However, for $j=1,2,\dotsc$, since $\omega_{n_{j}}\in K_{i}$ for $i\geq j+1$, so that $\lvert a_{i}\rvert(\omega_{n_{j}})\leq 2^{-i}$, we have
$a(\omega_{n_{j}})\geq a_{j}(\omega_{n_{j}})-\sum_{i=1}^{j-1}\lvert a_{i}\rvert(\omega_{n_{j}})-\sum_{i=j+1}^{\infty}\lvert a_{i}\rvert(\omega_{n_{j}})\geq j-2^{-j}.$
This shows that $\lim_{j\to\infty}a(\omega_{j})=\infty$.
Remark 2.6 and Remark 3.3 show that $\Omega$ is metrisable. To prove that $\Omega$ is complete with respect to $\mathcal{A}$, it suffices to show that Cauchy sequences in $\Omega$ are convergent. There is no need to consider general Cauchy nets. Lemma 5.8 shows that any Cauchy sequence has an accumulation point, and therefore it is convergent in $\Omega$. Thus Lemma 5.8 proves that (i) of Theorem 1.11 implies completeness of $\Omega$.
5.2 Proof of Theorem 1.18
@@
Proof 5.11.
As $K_{1},K_{2}$ are compact, and $\mathcal{F}$-convex we see that for $i=1,2$
@@
$\Big{\{}\omega\in\Omega\mid f(\omega)>\sup f(K_{2})\Big{\}},\text{ for all }f\in\mathcal{F},$ (5.4)
@@
$\Big{\{}\omega\in K_{4}\mid\frac{f_{i}(\omega)-\sup f_{i}(K_{1})}{\sup f_{i}(K_{2})-\sup f_{i}(K_{1})}\leq 1\text{ for }i=1,\dotsc,k\Big{\}}\subset\mathrm{int}K_{3}.$ (5.5)
@@
$p=\max\Big{\{}\frac{f_{i}-\sup f_{i}(K_{1})}{\sup f_{i}(K_{2})-\sup f_{i}(K_{1})}\mid i=1,\dotsc,k\Big{\}}\vee 0$ (5.6)
satisfies our requirements.
@@
$p\geq i\text{ on }\Omega\setminus\mathrm{int}K_{i+2}.$ (5.8)
@@
By continuity of $p$ we see that preimages of compacta are compact, i.e., $p$ is proper.
This completes the proof.
Proof 5.14 (Proof of Theorem 1.18).